OpenStudy (anonymous):

x^3+⅛ Need help factoring this please

OpenStudy (kc_kennylau):

you mean \(x^3+\dfrac12\)?!

OpenStudy (kainui):

That's 1/8.

OpenStudy (kc_kennylau):

wat?!

OpenStudy (anonymous):

\[x ^{3}+\frac{ 1 }{ 8 }\]

OpenStudy (kc_kennylau):

oh

OpenStudy (kc_kennylau):

Hint: \(\dfrac18=\left(\dfrac12\right)^3\) :)

OpenStudy (anonymous):

So \[(x+\frac{ 1 }{ 2 })(x+\frac{ 1 }{ 2 })(x+\frac{ 1 }{ 2 })\]

OpenStudy (kc_kennylau):

no... Do you recall the identity of the sum of cubes? :)

OpenStudy (kc_kennylau):

(In case you do not, \(a^3+b^3=(a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

Oh okay, i think i got it now, thank you! :)

OpenStudy (kc_kennylau):

no problem :)